The sum of the speeds of two trains is 723.6 miles per hour. If the speed of the first train is 10.4 mph faster than the second train, find the speeds of each.
Answer by rfer(16322)

x+x+10.4=723.6
2x=723.6-10.4
2x=713.2
x=713.2/2
x=356.6
x+10.4=367